OpenStudy (anonymous):

I cant seem to get the answer 44 when I am doing this question!!: Evaluate x^2y−2xy−y^2 when x =−3 and y=4. Steps please :)

OpenStudy (anonymous):

By hand not with a calculator

OpenStudy (anonymous):

\[x^2y-2xy-y^2\] everywhere you see x, put a -3, and everywhere you see y, put 4: \[(-3)^2(4)-2(-3)(4)-(4)^2 = (9)(4)+24-16 = 36+24-16 = 44\] If you need any steps explained just ask.

OpenStudy (anonymous):

Thank you so much! :)